Please show all calculations, no handwritten solutions

John Q. Investor takes a short position on 3,000 shares of Macy's(M), selling at $178.77 per share on February 11. Dividends are estimated at $24 per year paid quarterly, and the last Ex-date was December 31. JQ's broker charges him $500 for the loan of the shares every six months.JQ covers his position on May 12 at $188.91

a. At the time of the initial sell JQ's broker must deposit to JQ's margin account a total of $

b. and JQ's initial investment is

c. JQ's profit on this position is

d. JQ's holding period rate of return is

e. The rate of return on the underlying asset is
Answer #1

Short position of 3,000 shares selling at $178.77 each; Dividends = $24 per share

February 11, JQ will earn $178.77 * 3000 = $536,310

February 11, JQ will have to pay his broker $500

March 31, Dividend of $24 * 3000 = $72,000

May 12, JQ buys the stock for $188.91 * 3000 = $566,730

a) At time of initial sell, JQ's broker will deposit $536,310 + $500 = $536,810

This $500 will be added because JQ will have to deposit that amount in order to short sell the shares

b) JQ's initial investment is $500

c) JQ's profit position = ($536,310 + $72,000) - ($500 + $566,730) = $608,310 - $567,230 = $41,080

d) Holding period return = profit / investment = $41,080/$500 = 8216%

e) Rate of return of the underlying asset = (188.91 + 24-178.77)/178.77 = 19.1%